OpenStudy (anonymous):

nadia used the following process to find the y-intercept of the line described by the equation 2y+3x=18. step 1: subtract 3x from each side. 2y=18-3x. step 2: divide each side by 2. y=9-(3/2)x step 3: the y-intercept of y=mx+b is b. y-intercept is 9. according to nadia's method, which expression gives the y-intercept of the line described by ax+by=c?

OpenStudy (anonymous):

a) a/b b) -a/b c) c/b d) -c/b

OpenStudy (mathlegend):

c) c/b that is the y-intercept.

OpenStudy (mathlegend):

by+ax=c -ax -ax ------------ by = -ax+c (divide by b on both sides)) y= -ax+c/b